# Multiclass Local Calibration With the Jensen-Shannon Distance

### v1

@@ 8.1 Experimental Results
Q1: LoCal Nets achieve competitive global calibration results across datasets.
Figure 2 summarizes the performance of all methods on global calibration metrics across the three datasets. Our aim here is to show that LCN performs at a comparable level with the existing baselines. For $ECE$, DC is the strongest performer over cifar10 with $ECE\approx.0037\pm.0001$, followed by LCN with $ECE\approx.0047\pm.0003$ (Figure 2(a)). For $ECCE$, DC again leads with $\approx.0008\pm.0001$, while LCN achieves a level comparable to the other baselines. On cifar100, our method (LCN) achieves the best performance in terms of $ECE$ and is tied with DC and IR on $ECCE$ (Figure 2(b)). On tissuemnist, DC again performs best with $ECE=\approx.0044\pm.0002$, while LCN ranks second (Figure 2(c)). For $ECCE$, LCN performs slightly worse, obtaining $\approx.005\pm.001$.
In summary, DC is the overall best-performing method for global calibration, with its most pronounced advantage on balanced, low-class datasets such as cifar10. Still, our proposed LCN consistently achieves competitive results across datasets.
Q2: LoCal Nets consistently achieve superior results on local calibration metrics.
@@
Figure 3 evaluates the methods in terms of local calibration. Across all datasets, LCN consistently emerges as the best-performing approach, providing substantial improvements over competing methods. On cifar10, LCN achieves the lowest $LCE$ ($0078\pm.0002$), followed by DC ($010\pm.0003$), with TS and IR performing at a similar level to DC (Figure 3(a)). In terms of $MLCE$, LCN significantly outperforms all competitors with differences as large as .17.17. On cifar100, LCN again attains the best $LCE$ setting an evident gap with all the competitors. (Figure 3(b)). Here, PS achieves a $MLCE$ of $7030\pm.0366$, nearly matching LCN ($7022\pm.0063$). However, the lower variance of LCN highlights its reliability. The relative competitiveness of PS on this dataset is due to the limited per-class sample size (at most 600 instances), which favors parametric methods such as PS, whereas LCN benefits more from larger sample sizes. The largest relative gains are observed on tissuemnist (Figure 3(c)), where LCN achieves $00143\pm.0011$, markedly outperforming DC ($0028\pm.0011$) and all other baselines.
@@
Overall, these results demonstrate that LCN provides significant improvements in local calibration, particularly in settings with sufficient sample size.
Q3: Beyond calibration, LoCal Nets enhance predictive performance.
@@
Table 1 reports results on predictive performance, measured by $NLL$ and $Acc$. Our method achieves the largest reductions in $NLL$ across datasets, with the sole exception of cifar10, where LCNs rank second ($347\pm.002$) compared to DC ($333\pm.007$). On cifar100, LCN clearly outperforms all competitors with the lowest $NLL$ ($1.125\pm.002$ vs. $1.265\pm.016$ for DC). A similar trend is observed on tissuemnist, where LCN achieves $1.012\pm.003$, again surpassing DC ($1.052\pm.009$). Notably, LCN is the only method that improves model accuracy across datasets: $0.4\%$ on cifar10, $1.9\%$ on cifar100, and $2.7\%$ on tissuemnist. This occurs as other methods are forced to work only on logits, while LCN can learn new feature representations, possibly improving the predictions’ quality.
@@
9 Conclusions
In this work, we introduced a formal definition of multiclass local calibration and theoretically analyzed widely used calibration metrics under this assumption. Building on these insights, we proposed a novel post-hoc calibration method for neural networks that explicitly targets local calibration properties. Our empirical evaluation on both benchmarking and real-world datasets demonstrates that the proposed approach yields significant improvements in local calibration, while maintaining competitive performance with respect to global calibration metrics. These results highlight the importance of incorporating locality into the design of calibration methods.
Limitations and Future Work.
Although Theorem 4 establishes the consistency of our loss function, this guarantee holds only asymptotically. In finite-sample regimes, kernel estimates may suffer from non-negligible bias, which can limit performance. An important research direction is to explore adaptive kernel choices and scalable training procedures that better manage the bias–variance tradeoff inherent to local estimation. Moreover, the hyperparameter $\gamma$ requires careful tuning: small values yield unreliable estimates due to data sparsity, while large values obscure locality. A promising extension is to replace fixed kernels with adaptive or learned similarity functions, which may improve local calibration in high-dimensional or heterogeneous feature spaces. Finally, while our method is tailored to neural networks, extending it to enforce local calibration across other classes of models remains an important open direction.
Method Cifar10 Cifar100 TissueMNIST
AccAcc NLLNLL AccAcc NLLNLL AccAcc NLLNLL
LCN (Ours) $\mathbf{.888\pm.002}$ $\underline{.347\pm.002}$ $\mathbf{.688\pm.001}$ $\mathbf{1.125\pm.002}$ $\mathbf{.630\pm.001}$ $\mathbf{1.012\pm.003}$
DC $884\pm.001 .$$\mathbf{.333\pm.007}$ $670\pm.002$ $\underline{1.265\pm.016}$ $603\pm.008$ $\underline{1.052\pm.009}$
IR $884\pm.001 .$$364\pm.008 .$$670\pm.002$ $1.437\pm.029$ $603\pm.008$ $1.096\pm.016$
PS $884\pm.001 .$$466\pm.004 .$$670\pm.002$ $1.618\pm.007$ $603\pm.008$ $1.180\pm.008$
TS $884\pm.001 .$$362\pm.008 .$$670\pm.002$ $1.277\pm.009$ $603\pm.008$ $1.112\pm.023$
@@ Acknowledgements
The research of Cesare Barbera, Giovanni De Toni, Andrea Passerini, and Andrea Pugnana was partially supported by the following projects: Horizon Europe Programme, grants #101120237-ELIAS and #101120763-TANGO. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. The research of Cesare Barbera, Giovanni De Toni, Andrea Passerini, and Andrea Pugnana was also supported by Ministero delle Imprese e del Made in Italy (IPCEI Cloud DM 27 giugno 2022 – IPCEI-CL-0000007), PNRR-M4C2-Investimento 1.3, Partenariato Esteso PE00000013-“FAIR-Future Artificial Intelligence Research”, funded by the European Commission under the NextGeneration EU programme, and SoBigData.it, receiving funding from European Union – NextGenerationEU – National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) – Project: “SoBigData.it – Strengthening the Italian RI for Social Mining and Big Data Analytics” – Prot. IR0000013 – Avviso n. 3264 del 28/12/2021.
@@ Appendix A General Assumptions
Throughout this appendix, we present the theoretical results and corresponding proofs under a unified framework that simultaneously covers multi-dimensional binning-based calibration metrics and class-wise binning-based metrics constructed using equal-frequency binning. This formulation is adopted for clarity and compactness, as it allows for a seamless integration of both frameworks under a shared notation. For completeness, at the end of each proof that requires it, we explicitly discuss how the same analytical results extend to the case of class-wise calibration metrics based on fixed-confidence-threshold binning, which involves only minor technical adjustments. Hence, the presented analysis extends naturally to both binning schemes.
@@
  • Conditioning on kernel estimates. We condition on the kernel estimates $\hat{\theta}_{c}(\mathbf{y}_{i}\mid\mathbf{x}_{i})$, which are computed on an independent, unlimited disjoint dataset. For the evaluation set, these estimates depend only on the observed $\mathbf{x}_{i}$; thus, once features are fixed, the estimates are deterministic as well.
@@ A.3 Kernel Consistency Assumptions
It is important to acknowledge that estimates do not constitute a point-wise unbiased approximation of the true conditional distribution. Kernel estimators are known to suffer from both design bias—a form of bias introduced by the distribution of the covariates xx—and boundary bias, which is particularly pronounced near the edges of the support or in low-density regions of the input space [26]. Nevertheless, NW estimators are known to be consistent in the mean squared error (MSE) sense under mild conditions on the kernel and bandwidth sequence [8, 26]. Note that the MSE consistency is a stronger notion that implies Mean Integrated Squared Error (MISE) and Mean Absolute Error (MAE) consistencies. As a consequence, the estimator asymptotically yields a statistically meaningful approximation of the target function:
@@ B.1 Proof of Theorem 1
Given a locally calibrated classifier $f$, meaning that for instance $i\in D$, $\left\|f(\mathbf{x}_{i})-\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})\right\|_{1}\leq\varepsilon,$ where $\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})$ is a kernel estimator of the true conditional expectation $\mathbb{E}[\mathbf{y}_{i}\mid\mathbf{x}_{i}]$, we aim to bound the Multidimensional Expected Calibration Error (MECE), defined as:
$\mathbb{E}_{x\sim p(x)}\mathbb{E}_{y\sim p(y)}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right],$
by showing there exists $k\in[\nicefrac{{1}}{{C}},1]$ such that the following holds:
$MECE\leq\varepsilon\cdot k.$
@@
$\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\mathbb{E}_{y\sim p(y)}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right]=\int_{\mathcal{X}}\left[\sum_{y\in\mathcal{Y}}p(y)\cdot\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right]p(\mathbf{x})\,d\mathbf{x}$
$\leq\int_{\mathcal{X}}\left[\sum_{y\in\mathcal{Y}}\max_{y\in\mathcal{Y}}p(y)\cdot\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right]p(\mathbf{x})\,d\mathbf{x}$
@@
$\mathbb{E}_{x\sim p(x)}\mathbb{E}_{y\sim p(y)}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right]\leq\max_{y\in\mathcal{Y}}{{p(y)}}\int_{\mathcal{X}}\Bigg(\left\|\hat{\theta}(\mathbf{y}\mid\mathbf{x})-f(\mathbf{x})\right\|_{1}+\left\|{\mathbb{E}}[\mathbf{y}\mid\mathbf{x}]-\hat{\theta}(\mathbf{y}\mid\mathbf{x})\right\|_{1}\Bigg)\,p(x)\,dx.$
@@
$\mathbb{E}_{x\sim p(x)}\mathbb{E}_{\mathbf{y}\sim p(y)}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right]\leq(\varepsilon+\delta_{n})\cdot k.$
@@
$\mathbb{E}_{x\sim p(x)}\mathbb{E}_{y\sim p(y)}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(x)\right|\right]\leq\varepsilon\cdot k.\qed$ (11)
B.1.1 Proof of Theorem 2
In the following we prove that the value of a calibration metric satisfying the requisites of Definition 2 for a model that satisfies local calibration is bounded. More precisely, the proof will show that for a generic bin used to compute the metric, the calibration error in the bin is bounded. Finally, we will extend this to all bins simultaneously to bound the calibration error captured by the metric for a model that exhibits local calibration.
Let a deterministic binning function $\beta:\Delta_{c}\rightarrow\{1,\ldots,m_{B}\}$ partition the probability simplex $\Delta^{C}$ into $B$ disjoint bins $\{B_{b}\}_{b=1}^{m_{B}}$. For each bin $B_{b}$, define the index set of points that fall into it as $I_{b}=\{i:\hat{\mathbf{p}}_{i}\in B_{b}\}$, and let $|B_{b}|$ denote its cardinality. Define the per-bin frequency and confidence:
$\text{freq}_{b,c}=\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\mathbf{1}\{y_{i}=c\},\qquad\text{conf}_{b,c}=\frac{1}{|B_{b}|}\sum_{i\in I_{b}}f_{c}(\mathbf{x}_{i}),$
Let $\varphi:[0,1]\times[0,1]\rightarrow\mathbb{R}_{\geq 0}$ a scalar comparator that measure discrepancy between an empirical frequency and a predicted confidence that is Lipschitz in both arguments:
$\left|\varphi(a,b)-\varphi(a^{\prime},b^{\prime})\right|\leq L_{\varphi}(\left|a-a^{\prime}\right|+\left|b-b^{\prime}\right|)\quad\forall a,a^{\prime},b,b^{\prime}\in[0,1]$
The general multiclass bin-based calibration error metric is obtained as:
$\mathcal{E}(D;\varphi;\beta)=\sum_{b=1}^{m_{B}}w_{b}\sum_{c=1}^{C}\pi_{c}\cdot\varphi(\text{freq}_{b,c},\text{conf}_{b,c}).$ (12)
where $w_{b}$ are deterministic bin class weights typically set to $|B_{b}|/n$ and $\pi_{c}$ are deterministic class weights typically set to $1/C$ for balanced data. We now prove that if the model satisfies local calibration, i.e.,
$\left\|f(\mathbf{x}_{i})-\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})\right\|_{1}\leq\varepsilon,\quad\forall i\in\{1,\dots,n\},$
then, with probability at least as high as $1-\delta\in[0,1]$, the metric is bounded as follows:
$\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\Bigg[\varepsilon+\sum_{b=1}^{m_{B}}w_{b}\sqrt{\frac{\log(2Cm_{B}/\delta)}{2|\Psi(b,\mathcal{Y})|}}\Bigg].$
Where $\Psi(\cdot;\cdot)$ is a function that selects a bin based on index $b$ and labels $\mathcal{Y}$.

### v2

@@ 8.2 Experimental Results
Q1: Local Nets consistently achieve superior results on local calibration metrics. Figure 2 evaluates the methods in terms of local calibration. Across all datasets, LN consistently emerges as the best-performing approach, providing substantial improvements over competing methods.
On cifar10, LN achieves the lowest $LCE$ ($0079\pm.0003$), followed by KO ($0086\pm.0003$), with KC performing at a similar level to KO. In terms of $MLCE$, LN significantly outperforms all competitors with differences as large as $\approx.16$.
@@
On cifar100, LN again attains the best $LCE$, setting a clear gap with all competitors. Here, LN achieves $MLCE\approx.7022\pm.0070$, with PS nearly matching its result $(MLCE\approx.7031\pm.0409)$. However, the lower variance of LN highlights its reliability. The relative competitiveness of PS on this dataset is due to the limited per-class sample size (at most 600 instances), which favors parametric methods such as PS, whereas LN benefits more from larger sample sizes.
Notably, LN achieves the best performance in terms of $LCE$ also on tissuemnist, with $LCE\approx.0144\pm.0012$, outperforming KC ($LCE\approx.0165\pm.0008$) and all other baselines. In terms of $MLCE$, KO achieves the best performance with $6913\pm.0349$ while LN reaches similar results $(MLCE\approx.7293\pm.0359)$.
@@
Overall, these results demonstrate that LN provides significant improvements in local calibration, particularly in settings with sufficient sample size.
Q2: Local Nets achieve competitive global calibration results across datasets. Figure 3 summarizes the performance of all methods on global calibration metrics. We aim to show that LN performs at a level comparable to the existing baselines.
For $ECE$, DC is the strongest performer over cifar10 with $ECE\approx.0037\pm.0001$, followed by KO with $ECE\approx.0045\pm.0002$. Still, LN reaches comparable performance $(ECE\approx.0048\pm.0003)$. For $ECCE$, DC again leads with $\approx.0011\pm.0001$, while LN achieves a level comparable to the other baselines. On cifar100, LN achieves the best performance in terms of $ECE$ and is tied with DC and IR on $ECCE$ (Figure 3). On tissuemnist, DC again performs best with $ECE\approx.0062\pm.0007$, while LN ranks third (Figure 3). For $ECCE$, LN performs slightly worse, obtaining $\approx.0050\pm.0011$.
In summary, DC is the overall best-performing method for global calibration, with its most pronounced advantage on balanced, low-class datasets such as cifar10. Still, our proposed LN consistently achieves competitive results across datasets.
@@
Q3: Beyond calibration, Local Nets enhance predictive performance. Table 1 reports results on predictive performance, measured by $NLL$ and $ACC$. Our method achieves the largest reductions in $NLL$ across datasets, with the sole exception of cifar10, where LNs rank second ($347\pm.002$) compared to DC ($332\pm.007$). On cifar100, LN clearly outperforms all competitors with the lowest $NLL$ ($1.125\pm.002$ vs. $1.265\pm.016$ for DC). A similar trend is observed on tissuemnist, where LN achieves $1.012\pm.003$, surpassing KC ($1.021\pm.002$).
Like all the non-accuracy-preserving methods, we observe gains for LN: $0.4\%$ on cifar10, $1.9\%$ on cifar100, and $2.7\%$ on tissuemnist. This occurs as other methods are forced to work only on logits, while LN can learn new feature representations, possibly improving the predictions’ quality.
@@
9 CONCLUSIONS AND FUTURE WORK
Conclusion. In this work, we introduced a formal definition of multiclass local calibration and theoretically analyzed widely used calibration metrics under this assumption. Building on these insights, we proposed a novel post-hoc calibration method for neural networks that explicitly targets local calibration properties. Our empirical evaluation on both benchmarking and real-world datasets demonstrates that the proposed approach yields significant improvements in local calibration, while maintaining competitive performance with respect to global calibration metrics. These results highlight the importance of incorporating locality into the design of calibration methods.
Limitations and Future Work. Although Theorem 4 establishes the consistency of our loss function, this guarantee holds only asymptotically. In finite-sample regimes, kernel estimates may suffer from non-negligible bias, which can limit performance. An important research direction is to explore adaptive kernel choices and scalable training procedures that better manage the bias–variance tradeoff inherent to local estimation. Moreover, the hyperparameter $\gamma$ requires careful tuning: small values yield unreliable estimates due to data sparsity, while large values obscure locality. A promising extension is to replace fixed kernels with adaptive or learned similarity functions, potentially improving local calibration. Finally, while our method is tailored to neural networks, extending it to enforce local calibration across other classes of models remains an important open direction.
@@ Acknowledgements
We thank the anonymous reviewers for their feedback.
The research of Cesare Barbera, Giovanni De Toni, Andrea Passerini, and Andrea Pugnana was partially supported by the following projects: Horizon Europe Programme, grants #101120237-ELIAS and #101120763-TANGO. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. The research of Cesare Barbera, Giovanni De Toni, Andrea Passerini, and Andrea Pugnana was also supported by Ministero delle Imprese e del Made in Italy (IPCEI Cloud DM 27 giugno 2022 – IPCEI-CL-0000007), PNRR-M4C2-Investimento 1.3, Partenariato Esteso PE00000013-“FAIR-Future Artificial Intelligence Research”, funded by the European Commission under the NextGeneration EU programme.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] We clearly state the settings and assumptions of our method in Section 7.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Not Applicable] We do not study these properties in our work.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes] The code to reproduce our results can be found at https://anonymous.4open.science/r/local-calibration-25A3.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes] All our theorems and definitions in Sections 4, 5, 6, 7 clearly state assumptions.
    2. (b)
      Complete proofs of all theoretical results. [Yes] The proofs can be found in Appendix (Section Proofs)
    3. (c)
      Clear explanations of any assumptions. [Yes] See Sections 4, 5, 6, 7.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes] The code to reproduce our results can be found at https://anonymous.4open.science/r/local-calibration-25A3.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes] See Section 8.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes] We provide results using boxplots and $avg\pm std$.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes] We provide this information in Appendix (Section Experimental Details).
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes] We cite all the datasets owners.
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Appendix A General Assumptions
Throughout this appendix, we present the theoretical results and corresponding proofs under a unified framework that covers multi-dimensional binning-based calibration metrics This formulation is adopted for clarity and compactness. For completeness, at the end of each proof that requires it, we explicitly discuss how the same analytical results extend to the case of class-wise calibration metrics, which involves only minor technical adjustments. Hence, the presented analysis extends naturally to both binning schemes.
@@
  • Conditioning on kernel estimates. We condition on the kernel estimates $\hat{\theta}_{c}(\mathbf{y}_{i}\mid\mathbf{x}_{i})$, which are computed on an independent, arbitrarily large disjoint dataset. For the evaluation set, these estimates depend only on the observed $\mathbf{x}_{i}$; thus, once features are fixed, the estimates are deterministic as well.
@@ A.3 Kernel Consistency Assumptions
It is important to acknowledge that estimates do not constitute a point-wise unbiased approximation of the true conditional distribution. Kernel estimators are known to suffer from both design bias—a form of bias introduced by the distribution of the covariates xx—and boundary bias, which is particularly pronounced near the edges of the support or in low-density regions of the input space (Wasserman, 2006). Nevertheless, NW estimators are known to be consistent in the mean squared error (MSE) sense under mild conditions on the kernel and bandwidth sequence (Györfi et al., 2002; Wasserman, 2006). Note that the MSE consistency is a stronger notion that implies Mean Integrated Squared Error (MISE) and Mean Absolute Error (MAE) consistencies. As a consequence, the estimator asymptotically yields a statistically meaningful approximation of the target function:
@@ B.1 Proof of Theorem 1
Statement. Define the continuous Multidimensional Expected Calibration Error (MECE) as:
$M(f)=\sum_{y\in\mathcal{Y}}p(y)\,\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right],$
If a model $f$ satisfies local calibration, under standard regularity conditions there exists $k\in(0,1]$ such that continuous $MECE$ is asymptotically upper bounded:
$M(f)\leq\varepsilon\cdot k$ (12)
@@
$\sum_{y\in\mathcal{Y}}p(y)\,\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right]=\int_{\mathcal{X}}\left[\sum_{y\in\mathcal{Y}}p(y)\cdot\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right]p(\mathbf{x})\,d\mathbf{x}$
$\leq\int_{\mathcal{X}}\left[\sum_{y\in\mathcal{Y}}\max_{y^{\prime}\in\mathcal{Y}}p(y^{\prime})\cdot\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right]p(\mathbf{x})\,d\mathbf{x}$
@@
$\sum_{y\in\mathcal{Y}}p(y)\,\mathbb{E}_{x\sim p(x)}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right]\leq\max_{y\in\mathcal{Y}}{{p(y)}}\int_{\mathcal{X}}\Bigg(\left\|\hat{\theta}(\mathbf{y}\mid\mathbf{x})-f(\mathbf{x})\right\|_{1}+\left\|{\mathbb{E}}[\mathbf{y}\mid\mathbf{x}]-\hat{\theta}(\mathbf{y}\mid\mathbf{x})\right\|_{1}\Bigg)\,p(x)\,dx.$
@@
$\sum_{y\in\mathcal{Y}}p(y)\,\mathbb{E}_{x\sim p(x)}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right]\leq(\varepsilon+\delta_{n})\cdot k.$
@@
$\sum_{y\in\mathcal{Y}}p(y)\,\mathbb{E}_{x\sim p(x)}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(x)\right|\right]\leq\varepsilon\cdot k.\qed$ (13)
B.2 Proof of Theorem 2
Statement. Let $\mathcal{E}(D;\varphi;\beta)$ be a calibration metric satisfying Definition 2, evaluated on a sample $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$. Let $f$ be locally calibrated with error $\varepsilon$ relative to kernel estimates $\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})$ computed on an independent arbitrarily large auxiliary dataset. Conditioning on the evaluation inputs $\{\mathbf{x}_{i}\}_{i=1}^{n}$ and corresponding consistent kernel estimates, for any $\delta\in(0,1)$, with probability at least $1-\delta$:
$\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\sum_{b=1}^{m_{B}}\sum_{c=1}^{C}w_{b,c}\,\cdot\,\pi_{c}\Bigg(\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2|\Psi(b,c)|}}\;+\;\varepsilon\Bigg)$ (14)
where $\Psi(\cdot,\cdot)$ secelts a bin based on the index $b$ and possibly the label $c$.